Project

General

Profile

10/23/2013 Minutes

Attendance

Nick, Jan, Michael, Mat, Susan, Kanika, Evan, Dominick (Ok with Dom), Gavin

Operations

  • End of shift form now has has run synopsis
  • dcm-2-02-09: (problems) being swapped (11474 first untangled 2 diblock cold run)
    1. Diblock 1 is good
    2. Beam is off all day

Production

  • Cosmic trigger haven’t been processed for 1 week
    1. Nick waiting for NDOS to come through
    2. Gavin says nova pro can get a bogged down queue (can we stagger?)
    3. Copy latest file with priority

Outstanding Issues:

  • cosmic trigger crash issue
    1. major fix in the working to CAF Floating Point Exception
    2. We need a scheme to get this rolling (Branches)